of protecion. Poly and Digit become worrned thut their pets may not be covered by the Wizards protection spelil which is based castle to rescue their pets. Gourmandos storm is building to a fury, and the sices are he Kingelorms creatures one more time to see if they need to rush back to thn Question 3 of 4 Refer to the previous question answer to the previous question, as wellas the fact that Poly and Digts combined data has a mean of previous question, as well as the fact that Poly and Digits combined data has a mean of 5.9 and a standard deviation of 1.21 Which of the following statements is/are justified based on the data? There may be more than one correct answer. Select all that apply Approximately Approximately 68% of the heghts of people in the Kingdomare between 1.21 and 59 feet Approximately 95% of the heights of people in the Kingdom are between 59 and 832 feet Approximatey 68% of the-eights of peopie Approximately 95% orne rera of people in the Kingdom are between 3.48 and 8.32 fee. Approximatey 99.7%of tha hegrts of people i 99.7% of the heights of people in the Kingdom are between 2.27 and 953 feet. m the Kingcom are between 4.00 and 711 feet me Kirgdomare betr,een 2.27 ang 7.11%ee
usly to begin forming his spell of protection Poly and Digit become worried that their pets may not be covered by the Wizards protection spel oms creatures one more time to see iFthey need to rush back to the castle to rescue their pets Gourmandos storm is building to a fury, and t Question 4 of 4 Refer to the previous question Poly is anxious about her pet dragon, Lutwidge. who is 10 feet tall. Digt is nervous about his pet dragon. Charlemagne, who is 4 feet tall The Wizards spell is less effective for creatures whose heights are more than 2 standard deviations from the mean. Using be worried about their pets the 68 95-99.7 rule and your answers) from question 3, explain why Poly and Digit should or should not EX x,x五 Normal : BlyG
Answer #1

Please read the answer below. Don't hesitate to give a "thumbs up" in case you're satisfied with the answer.

3.

mean = 5.9
stdev = 1.21

within 1st dev ( which covers 68% of area) = 5.9+/- 1*1.21 = 4.69 to 7.11
within 2nd dev ( which covers 95% of area)= 5.9+/- 2*1.21 = 3.48 to 8.32
within 3st dev ( which covers 99.7% of area)= 5.9+/- 1*1.21 = 2.27 to 9.53

So, options 1st , 5th and 6th options is correct.